Question 9 (1 point) The stock price for International Business Machines (IBM) historically has followed an approximately normal distribution (when adjusting for inflation) with a mean of $166.305 and standard deviation of $4.2216. What is the probability that on a selected day the stock price is below $162.5? 1) We do not have enough information to calculate the value. 2) 0.8163 3) 0.1837 4) 0.1706 5) 0.8294
Answer #1

P(X < 162.5)

= P((X - \mu)/\sigma < (162.5 - \mu)/\sigma)

= P(Z < (162.5 - 166.305)/4.2216)

= P(Z < -0.901)

= 0.1837

Option - 3 is correct.